OpenStudy (anonymous):

Two planes which are 2265 miles a apart, fly toward each other. Their speed differs by 55 mph. If they pass each other in 3 hours, what is the speed of each?

OpenStudy (a_clan):

use relative velocity

OpenStudy (a_clan):

let their speeds be v1 and v2 Relative speed(or frame of reference); Assume one plane to be at rest, then the speed of other will become (v1+v2) v2-v1=55.......(1) Given so, v2=v1+55 so, (v1+v2)= v2+v2+55 = 2.v2+55 Now, distance/speed = time 2265/(2v2+55) = 3 hence v2 and v1

OpenStudy (a_clan):

v2=350 mph v1=295 mph